# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v2

@@ 4 Experiments
In this section, we evaluate the representations learned with our pretext tasks by measuring their ability to scale downstream performance with unlabelled data. This includes understanding how well they can generalise across datasets, subjects, and tasks. We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG [32] and faster sampling rates than fMRI [25].
@@ 4.1 Experimental setup
Datasets. Unless specified otherwise, our experiments use Cam-CAN [48, 53] as an unlabelled representation learning dataset for pre-training. This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings. For our downstream tasks, we use two labelled heard speech MEG datasets where participants listen to short stories or audiobooks. Armeni et al. [2] contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al. [24] has 27 subjects, each recorded for 2 hours (54 hours total). Overall, we utilise over 200 hours of data. To the best of our knowledge, this is the largest volume of MEG data ever used for speech decoding.
Preprocessing. Each recording is in $\mathbb{R}^{S\times T}$ where $S$ is the number of sensors and $T$ is the number of time points sampled by the scanner. To eliminate high-frequency muscle movement artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.5Hz to remove slow-drift artifacts. Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at multiples of 50Hz to account for line noise. Next, we downsample the signal to 250Hz, avoiding aliasing at frequencies up to our low-pass filter threshold. Finally, we detect bad sensor channels, those with significant noise and artifacts, using a variance threshold and replace them by interpolating the spatially nearest sensors.
@@
Training. We pre-train all models to completion and then shallow or deep fine-tune on labelled data for each task, using this to analyse the generalisability of the pre-trained representation. Appendix E provides complete training details for all experiments.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
We investigate whether our self-supervised objectives produce useful representations by first learning a representation with each pretext task independently and then analysing its downstream generalisation using shallow fine-tuning (i.e. with the representation frozen). In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, and critically, outperform a randomly initialised and shallow fine-tuned baseline. The results are statistically significant in all cases, providing initial evidence that all of our tasks are helpful in learning representations that generalise downstream to speech decoding.
Interestingly, the combination of all pretext tasks leads to better generalisation than any task on its own. As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding. For our other downstream task, voicing classification, we were not able to use shallow fine-tuning to generalise to a statistically significant degree with any of the pretext tasks nor with a random backbone. It may be too difficult to learn with shallow fine-tuning and so we turn to using deep fine-tuning for this task in the next section.
Pretext task Speech detection balanced accuracy
Armeni $t$ $p$ Gwilliams $t$ $p$
Amp($\rho=0.2$) $0.5846\pm 0.0032$ $26$ $7\mathrm{e}{-4}$ $0.5116\pm 0.0004$ $29$ $6\mathrm{e}{-4}$
Phase($\rho=0.5$) $0.5831\pm 0.0029$ $29$ $6\mathrm{e}{-4}$ $0.5093\pm 0.0014$ $7$ $1\mathrm{e}{-2}$
Band $0.5941\pm 0.0024$ $39$ $2\mathrm{e}{-3}$ $0.5177\pm 0.0028$ $6$ $1\mathrm{e}{-2}$
All $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
None (rand. backbone) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
Chance-level $0.5$- -- $0.5$- --
Table 1: Pre-training with pretext tasks produces generalisable representations. We provide the test accuracy at the point of best validation accuracy (early stopping). In the none baseline, the network backbone is randomly initialised before fine-tuning (i.e. there is no pre-training). When all pretext tasks are used, their losses are weighted equally. The uncertainty is the standard error of the mean over three random seeds. We quote the $t$-score and $p$-value for each result from a one-sample one-sided $t$-test where the population mean is chance-level accuracy. Models are pre-trained on all of Cam-CAN [48, 53] and fine-tuned on Armeni et al. [2] or Gwilliams et al. [24] respectively.
@@
Among the pretext tasks, band prediction performs best. This could be because the representation easily identifies phase-locking to speech onset in theta waves [43]. However, further investigation is necessary to determine which frequency band is most informative. The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix F). We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset [26]. In contrast, a larger $\rho$ for phase shift prediction may be best because phase is descriptive of neural synchrony which is distributed across the brain. Unlike localised responses, a large proportion of the sensors can detect this feature.
Whilst we have shown that our representations generalise, do they outperform supervised baselines when trained to saturation? Using shallow fine-tuning once more, we compare our approach directly to similarly parameterised supervised classifiers in Table 2. A representation pre-trained with all pretext tasks significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets. Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is far larger without an encoder. Even with this bias favouring the supervised classifier, the self-supervised case performs better. Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
Model Speech detection balanced accuracy
Armeni $t$ $p$ Gwilliams $t$ $p$
Supervised
Linear $0.5222\pm 0.0015$ $12$ $3\mathrm{e}{-3}$ $0.5006\pm 0.0001$ $5$ $2\mathrm{e}{-2}$
Self-supervised + fine-tuning
Random backbone
   MLP (two layers) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
Pre-trained backbone
   Linear $0.5853\pm 0.0017$ $52$ $3\mathrm{e}{-4}$ $0.5011\pm 0.0001$ $9$ $6\mathrm{e}{-3}$
   MLP (two layers) $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
Chance-level $0.5$- -- $0.5$- --
Table 2: Self-supervised pre-training outperforms supervised classifiers. We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion. We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds. We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance. Self-supervised models are pre-trained on all of Cam-CAN [48, 53] and shallow fine-tuned on Armeni et al. [2] or Gwilliams et al. [24] while the supervised models are trained entirely on the latter datasets.
@@ 4.3 Scaling Speech Decoding With Unlabelled Data
Figure 3: Scaling unlabelled data improves downstream generalisation. We pre-train the model on increasing amounts of unlabelled data from Cam-CAN [48, 53] and fine-tune on voicing and speech detection on the Armeni et al. [2] and Gwilliams et al. [24] datasets. No pre-training shows best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). We show only one no pre-training line as the result is the same for both datasets. The shaded area shows the standard error of the mean across three seeds.
Previously, we showed that the combination of all pretext tasks leads to the best representations and that this outperforms a randomly initialised backbone as well as a supervised classifier. In this section, we examine how these pretext tasks fare as we increase the amount of unlabelled data, analysing scaling performance on downstream tasks. As before, we pre-train with the combined pretext tasks using Cam-CAN [48, 53], which does not have any speech decoding labels, and now fine-tune for ten epochs on Armeni et al. [2] and Gwilliams et al. [24], showcasing transfer with minimal fine-tuning. This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI. Where earlier we analysed only speech detection, here, we also focus on voicing classification by using deep fine-tuning for this task only. As evidenced by Section 4.2, we require this because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
Figure 3 shows balanced accuracy as we increase the amount of unlabelled data in pre-training up to approximately 160 hours. For both tasks, pre-training outperforms the no pre-training baseline, which fails to generalise at all, showing that any unlabelled data is sufficient to pre-train a useful representation. In both tasks, there is a clear improvement in accuracy as the amount of unlabelled data increases. Thus, adding unlabelled data improves generalisation and scales performance. With speech detection, the improvement in accuracy is more significant for Armeni et al. [2]. Noting that there is more within-subject data in this dataset than Gwilliams et al. [24], speech detection may be more sensitive to subject variation than voicing classification.
@@
We scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our self-supervised method is an exception to the common consensus that pooling subjects worsens generalisation. Moreover, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is particularly notable as the Armeni et al. [2], Gwilliams et al. [24], and pre-training datasets all use different scanners entirely. Performing well across these datasets indicates that, together, our architecture and pretext tasks successfully generate representations that are generalisable across heterogeneous scanners.
Beyond hardware, our results also show improvements on distinct tasks, namely speech detection and voicing classification, thus illustrating cross-task generalisation. Thus, our pretext tasks are sufficiently generic to produce representations that generalise to multiple downstream tasks, indicating their applicability to speech decoding more generally. Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks. Remarkably, this shows that brain data collected from any task (including those that are not linguistic) can be used to scale downstream speech decoding performance.
4.4 Scaling Unlabelled Data Improves Zero-Shot Generalisation To Novel Subjects
@@
Brain data is variable across participants, leading to difficulty transferring models to novel subjects [11]. Whilst we have shown generalisation across subjects with our method, here, we investigate how well we can generalise to novel subjects—an even more difficult challenge. This is critical in order to widely deploy speech BCIs for new patients. In the following experiment, we pre-train using the same setup as Figure 3, but fine-tune only on Gwilliams et al. [24]. We use FiLM conditioning as this performs best for seen and unseen subjects (Appendix B). During training, we hold out three subjects with which we evaluate novel subject generalisation (i.e. unseen subject performance).
Figure 4: Scaling pre-training data improves novel subject generalisation. As before, we pre-train with increasing amounts of unlabelled data from Cam-CAN [48, 53], but now fine-tune only on Gwilliams et al. [24]. In the seen case, we evaluate on held-out sessions from subjects that were used in the training data. For the unseen case we evaluate on three completely held-out subjects. No pre-training shows the best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). We show only one no pre-training line as the result is the same for both datasets. The shaded area is the standard error of the mean over three seeds.
@@
Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive trend in performance for unseen (novel) subjects. This is a promising result, indicating that scaling with our method is an encouraging direction for resolving the issues related to individual variance faced by prior work.

### v3

@@ 4 Experiments
In this section, we evaluate the representations learned with our pretext tasks by measuring their ability to scale downstream performance with unlabelled data. This includes understanding how well they can generalise across datasets, subjects, and tasks. We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG (Lopes da Silva, 2013) and faster sampling rates than fMRI (Hall et al., 2014).
We pre-train all models to completion and then fine-tune on labelled data for each task. In all tables and figures, we quote the receiver operating characteristic area under the curve (ROC AUC) where chance is always 0.5 regardless of class balance. We show the test ROC AUC at the best validation ROC AUC (early stopping) and quote uncertainty as the standard error of the mean over three seeds. Additionally, we state the $t$-score and $p$-value from single-sample one-sided $t$-tests against chance.
@@ 4.1 Experimental setup
Datasets. Unless specified otherwise, our experiments use Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) as an unlabelled representation learning dataset for pre-training. This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings. For our downstream tasks, we use two labelled heard speech MEG datasets where participants listen to short stories or audiobooks. Armeni et al. (2022) contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al. (2023) has 27 subjects, each recorded for 2 hours (54 hours total). Overall, we utilise over 200 hours of data. To the best of our knowledge, this is the largest volume of MEG data ever used for speech decoding.
Preprocessing. Each recording is in $\mathbb{R}^{S\times T}$ where $S$ is the number of sensors and $T$ is the number of time points sampled by the scanner. To eliminate high-frequency muscle movement artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.5Hz to remove slow-drift artifacts. Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at multiples of 50Hz to account for line noise. Treating the low-pass filter threshold as the Nyquist frequency, we downsample the signal to twice that at 250Hz, avoiding aliasing within our band of interest. Finally, we detect bad sensor channels, those with significant noise and artifacts, using a variance threshold and replace them by interpolating the spatially nearest sensors.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
Our first experiment investigates whether our self-supervised objectives produce generalisable representations. In Table 1, we show the results of pre-training models with each pretext task independently as well as together. Here, all of our pretext tasks lead to results that are statistically significant, and outperform a baseline fine-tuned without pre-training. This provides initial evidence that our tasks are helpful in speech decoding. Interestingly, the combination of all pretext tasks leads to better generalisation than any task on its own. As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding than any individual task.
Armeni Gwilliams
Experiment ROC AUC $t$ $p$ ROC AUC $t$ $p$
Linear $0.559\pm 2\mathrm{e}{-4}$ $341$ $4\mathrm{e}{-6}$ $0.527\pm 7\mathrm{e}{-5}$ $379$ $3\mathrm{e}{-6}$
BIOT + linear $0.500\pm 4\mathrm{e}{-4}$ 00 $6\mathrm{e}{-1}$ $0.499\pm 2\mathrm{e}{-4}$ $-3$ $1\mathrm{e}{+0}$
Ours No pre-training $0.519\pm 0.002$ $8$ $7\mathrm{e}{-3}$ $0.498\pm 0.003$ 00 $7\mathrm{e}{-1}$
+ linear Amp($\rho=0.2$) $0.602\pm 0.001$ $114$ $4\mathrm{e}{-5}$ $0.532\pm 0.005$ $6$ $1\mathrm{e}{-2}$
Phase($\rho=0.5$) $0.603\pm 0.003$ $35$ $4\mathrm{e}{-4}$ $0.535\pm 0.003$ $12$ $3\mathrm{e}{-3}$
Band $0.616\pm 0.003$ $44$ $3\mathrm{e}{-4}$ $0.542\pm 0.001$ $46$ $2\mathrm{e}{-4}$
All tasks $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $\mathbf{0.543}\pm 0.003$ $13$ $3\mathrm{e}{-3}$
Table 1: Pre-training with pretext tasks leads to better representations for decoding speech. In the linear-only case, we train a supervised linear classifier on the input MEG signals. For BIOT, we train a linear layer on top of the pre-trained BIOT representation, with the rest of the model frozen. Similarly, for ours, we train a linear probe on top of our pre-trained backbone with its weights frozen. In the no pre-training baseline, the backbone uses randomly initialised and subsequently unmodified weights. When all pretext tasks are used, their losses are weighted equally.
Now, we turn to the other baselines. Our approach significantly outperforms the equivalent with a raw MEG input instead of a pre-trained representation (the linear experiment). Here, the baseline has substantially more trainable parameters because the input dimension is far larger without an encoder. Even with this bias favouring the experiment with the raw input, using our representation still performs better. We also compare our approach to BIOT (Yang et al., 2023) which is a similar state-of-the-art self-supervised method. When BIOT is pre-trained using exactly the same data, the fine-tuned probe fails to generalise entirely after exhaustive hyperparameter tuning. We put this down to three critical reasons. Firstly, BIOT was designed around considerably lower-dimensional signals. Their EEG evaluation used an order of magnitude fewer sensors than our MEG data. With MEG, their transformer approach requires many more channel embeddings, leading to difficulty learning the complex interactions between sensors. Secondly, our self-supervised objective extracts speech decoding features which is essential for solving speech decoding tasks. BIOT performs well on simple EEG tasks in Yang et al. (2023)’s evaluation, but non-invasive speech decoding is significantly more challenging. Together, these obstacles suggest a vast amount of data is required to learn their objective with MEG. Indeed, given that they pre-train with over 50 thousand hours of EEG data in their evaluation, their objective appears too general to efficiently learn a representation for speech decoding from the limited amount of MEG pre-training data (160 hours) available to us. This highlights the importance of data-efficiency in SSL methods for MEG.
@@
Among the individual pretext tasks, band prediction leads the rest. Perhaps this is because, by learning to discriminate between meaningful bands, the representation easily identifies phase-locking to speech onset in theta waves (Peelle et al., 2012). Further investigation is necessary here. The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search. We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset (Hamilton et al., 2018). In contrast, a larger $\rho$ for phase shift prediction could lead to representations that better discriminate neural synchrony information which is distributed across the brain rather than localised. As a result, a large proportion of the sensors in a MEG scanner should encode information about this feature.
@@ 4.3 Scaling Speech Decoding With Unlabelled Data
Figure 3: Scaling unlabelled data improves generalisation. We pre-train the model on increasing amounts of unlabelled data from Cam-CAN (Shafto et al., 2014; Taylor et al., 2017). The solid lines are the best linear fits to the data and the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
Here, we analyse generalisation as we increase the volume of unlabelled data, analysing scaling performance on downstream tasks. As before, we pre-train with the combined pretext tasks. Figure 3 shows ROC AUC as we increase the amount of unlabelled data in pre-training up to approximately 160 hours. For both tasks, pre-training with any amount of data is sufficient to beat chance and there is a clear improvement in accuracy as the amount of unlabelled data increases. For speech detection on Armeni et al. (2022), scaling appears logarithmic in log-space; for all others, ROC AUC improves log-linearly within the data regime we study. In any case, adding unlabelled data has improved generalisation. Notably, we have scaled far beyond the data regime of prior surgical and non-surgical work and yet performance has continued to scale. Thus, our self-supervision approach may remain useful as the volume of open data in the field continues to rapidly increase.
@@
Our results also reveal several new and notable phenomena. Firstly, we scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our method is an exception to the common consensus that pooling subjects worsens generalisation. Secondly, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is particularly surprising as the Armeni et al. (2022), Gwilliams et al. (2023), and our pre-training dataset all use different scanners entirely. Performing well across these datasets indicates that, together, our architecture and pretext tasks successfully generate representations that are generalisable across heterogeneous scanners. Finally, we note that our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks. Remarkably, this shows that unlabelled brain data collected from any task (including those that are not linguistic) can be used to improve speech decoding performance.
Since the results show improvements on both downstream tasks, this indicates that our pretext tasks are sufficiently generic to produce representations that work with multiple speech decoding tasks while still generalising well on each task individually. This is generally a challenging trade-off to manage. However, we notice that in both tasks, the base accuracy is higher and the improvement in ROC AUC is steeper for Armeni et al. (2022). This is likely to be because this dataset has more within-subject data. The weaker results for Gwilliams et al. (2023) may be a consequence of the larger number of subjects with shorter intra-subject recordings and greater subject variation. These observations support the findings of other recent work such as Csaky et al. (2022).
4.4 Scaling Unlabelled Data Improves Generalisation To Novel Subjects
@@
In neuroimaging, brain data is generally highly variable across participants, leading to difficulty transferring models to novel subjects (Csaky et al., 2022). Whilst we have shown generalisation across subjects, here, we investigate whether we can generalise to novel subjects—an even more difficult challenge. This is critical in order to widely deploy speech BCIs for new patients. In this experiment, we fine-tune only on Gwilliams et al. (2023) and hold out three subjects with which we evaluate novel subject generalisation.
Figure 4: Scaling unlabelled data improves novel subject generalisation. We fine-tune on Gwilliams et al. (2023). When in-distribution, we evaluate on held-out sessions from subjects in the training set; when out-of-distribution, we evaluate on three held-out subjects. The solid lines are the best linear fits while the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
@@
Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive log-linear trend in performance for novel subjects. This indicates that scaling our method is an encouraging direction for resolving the challenges of subject variance faced by prior work. Moreover, as far as we are aware, this is the first result to demonstrate novel subject generalisation in speech decoding from MEG.
4.5 Aggregating Unlabelled MEG Datasets
To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets. As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS (Schoffelen et al., 2019) and Cam-CAN (Shafto et al., 2014; Taylor et al., 2017). In this section, we investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
Pre-training Armeni Gwilliams
dataset Hours ROC AUC $t$ $p$ ROC AUC $t$ $p$
Cam-CAN 159 $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $0.543\pm 0.003$ $13$ $3\mathrm{e}{-3}$
MOUS 160 $0.605\pm 0.000$ $261$ $7\mathrm{e}{-6}$ $0.543\pm 0.004$ $9$ $5\mathrm{e}{-3}$
Cam-CAN + MOUS 319 $0.611\pm 0.003$ $40$ $3\mathrm{e}{-4}$ $\mathbf{0.546}\pm 0.002$ $20$ $1\mathrm{e}{-3}$
Table 2: Combining unlabelled datasets shows signs of outperforming single studies. We examine performance on the speech detection task. We see a small improvement when the datasets are combined on Gwilliams et al. (2023), but not Armeni et al. (2022).
The results in Table 2 show, for the first time, that combining datasets can improve performance on downstream speech decoding tasks. It leads to better performance on Gwilliams et al. (2023) compared to pre-training on either dataset alone. Interestingly, this was not the case for Armeni et al. (2022) where pre-training on Cam-CAN alone performed best. Combined pre-training did, however, outperform training only on MOUS. It is surprising that pre-training on Cam-CAN was better than pre-training on MOUS when evaluating on Armeni et al. (2022) given that MOUS and Armeni et al. (2022) both used speech tasks and were acquired on the same MEG scanner. Cam-CAN, by contrast, did not use a speech task and was acquired on a different MEG scanner. We hypothesise that the better results for Cam-CAN are due to it being a cleaner dataset. During our experiments, we found that data quality, even among unlabelled data, can have a significant affect as artefacts in recordings disrupt learning.
While the combination of the two datasets includes far more hours of data than any prior work on deep learning with MEG, further work needs to be done to aggregate more datasets. Here, we were limited by compute budget. Increasing the number of datasets could enable the network to eventually always improve over the best singular dataset. Just as increasing the number of subjects (rather than only within-subject data) improves novel subject generalisation, a larger number of datasets may be key to scaling results when datasets are aggregated in pre-training.